# Emergent World Models and Latent Variable Estimation in Chess-Playing Language Models

### v1

@@ 3.1 An Internal Board State Representation
This method involves training separate linear probes for each of the model’s layers and each individual square on the board, where each probe at layer takes the activations after layer as input to predict the state of square . We follow Nanda et al. (2023)’s finding that the model represents squares as containing (Mine, Yours, Empty) rather than (Black, White, Empty). To use this objective, we must train separate linear probes for predicting the board state at every white move and every black move, rather than a single linear probe for every move.
We train our probes on 10,000 games not found in the model’s training dataset. An additional set of 10,000 games, also excluded from the model’s training dataset, serves as our held out test set. As detailed in Table 1, the most accurate probe achieves a 99.6% accuracy in classifying the state of each square across the test games. Further insights into the model’s performance across different layers are detailed in Figure 3. Most notably, the 8 layer model obtains peak accuracy at layer 5, while the 16 layer model doesn’t obtain comparable accuracy until layer 11. Intuitively, one might expect that the most efficient approach would be calculating the board state as soon as possible and then using this information to calculate the next move. Instead, we see the deeper model using twice as many layers to calculate its ultimately more accurate board state.
We can visualize the model’s internal representation of the board by creating heat maps. These heat maps were derived from the probe outputs, which had been trained on a one-hot objective to predict whether a chess piece, such the black king, was present on a given square (1 if present, 0 if not). As seen in Figure 1, the model’s internal representation reflects reasonable facts about the board state. For example, the internal representation is very confident that no white pawns are on either back rank.
@@ 4.2 Board State Interventions
Our board state intervention is detailed in Figure 4. We test our intervention on 5,000 different board states. In our intervention, we delete a strategically relevant piece (the piece the model moved under the original board state) from the original board to obtain a modified board. We modify the model’s activations to reflect this new board state, then sample a new move from our modified model.
To evaluate our intervention, we sample five probabilistic moves from both the original and modified model at a temperature of 1.0. The original model provides a baseline legal move rate with no intervention applied, and the modified model’s legal move rate measures the success of the intervention. Our intervention significantly outperforms our baseline as detailed in Table 2. However, the best intervention success rate obtained is 92%, indicating that there is potential for improved model intervention strategies.
@@ 4.3 Model Skill Interventions
We obtain a skill vector by obtaining the difference vector between the high skill and low skill vectors derived from the skill probe. Alternatively, we can add the contrastive activation derived skill vector to the model’s activations. We can flip the sign of the intervention to decrease the model’s skill. An excessively large intervention will cause the model to output illegal moves and random characters. When testing our intervention to both increase and decrease model skill, we ensure that the model still makes over 98% legal moves after the intervention.
@@ 5 Related Work
There have been several examples of researchers finding semantics extraction in Large Language Models trained on text. In 2017, researchers trained a long short-term memory network (LSTM) to predict the next character in Amazon reviews found that linear probes trained on the model exceeded the state of the art in sentiment classification (Radford et al., 2017). More recently, researchers have found evidence of concepts such as truthfulness and emotion in the Llama LLM (Li et al., 2023b; Zou et al., 2023).
@@ 6 Conclusion
Chess is a constrained setting compared to the domain of natural language. This constrained setting makes it possible to perform these nuanced measurements and interventions and provide valuable insights into the internal workings of large language models. In addition, it also enables exploration and comparisons on a variety of probing and intervention methods. There is room for improvement on our intervention approaches, and we can easily measure our ability to make fine grained changes to the model’s activations, such as by deleting a single piece from the board.
However, chess is ultimately a well defined, rule based system that does not capture the complexity and ambiguity present in less structured environments like natural language. Accurate understanding of LLM world models could enable many applications, such as detecting and reducing hallucinations. Our hope is that these techniques can become practical in these domains.
@@ Appendix C Intervention Technique Comparisons
We find that a single layer intervention works almost as well as a multiple layer intervention. We had the following board intervention success rates per single layer intervention.

### v2

@@ 3.1 An Internal Board State Representation
This method involves training separate linear probes for each of the model’s layers and each individual square on the board, where each probe at layer takes the activations after layer as input to predict the state of square . We follow Nanda et al. (2023)’s finding that the model represents squares as containing (Mine, Yours, Empty) rather than (Black, White, Empty). To use this objective, for every layer, we must train one group of 64 linear probes for predicting the board state at every white move and a second group of 64 linear probes for every black move, rather than a single group of 64 linear probes for every move.
We train our probes on 10,000 games not found in the model’s training dataset. An additional set of 10,000 games, also excluded from the model’s training dataset, serves as our held out test set. As detailed in Table 1, the most accurate probe achieves a 99.6% accuracy in classifying the state of each square across the test games. Further insights into the model’s performance across different layers are detailed in Figure 3. Most notably, the 8 layer model obtains peak accuracy at layer 6, while the 16 layer model doesn’t obtain comparable accuracy until layer 12. Intuitively, one might expect that the most efficient approach would be calculating the board state as soon as possible and then using this information to calculate the next move. Instead, we see the deeper model using twice as many layers to calculate its ultimately more accurate board state.
We can visualize the model’s internal representation of the board by creating heat maps. These heat maps were derived from the probe outputs, which had been trained on a one-hot objective to predict whether a chess piece, such the black king, was present on a given square (1 if present, 0 if not). As seen in Figure 1, the model’s internal representation reflects reasonable facts about the board state. For example, the internal representation is very confident that the black king is not on the white side of the board.
@@ 4.2 Board State Interventions
Our board state intervention is detailed in Figure 4. We test our intervention on 5,000 different board states. Intuitively, we find a piece the model intends to move and remove the piece from the model’s internal board state by modifying the model’s activations (we do not alter the input PGN string), then sample additional moves from the model. If our intervention is successful, the model’s moves will be legal under the new hypothetical board state.
To evaluate our intervention, we sample five probabilistic moves from both the original and modified model at a temperature of 1.0. The original model provides a baseline legal move rate with no intervention applied, and the modified model’s legal move rate measures the success of the intervention. Our intervention significantly outperforms our baseline as detailed in Table 2. However, the best intervention success rate obtained is 92%, indicating that there is potential for improved model intervention strategies. As seen in Figure 5, the intervention can have unintended side effects and make the positions of other pieces less distinct.
@@ 4.3 Model Skill Interventions
We obtain one skill vector from contrastive activations and the second by subtracting the linear probe’s low skill vector from its high skill vector. We can add the skill vector to the model’s activations to increase its skill. We can flip the sign of the intervention to decrease the model’s skill. An excessively large intervention will cause the model to output illegal moves and random characters. When testing our intervention to both increase and decrease model skill, we ensure that the model still makes over 98% legal moves after the intervention. Although the skill probe was trained to predict the Elo ratings of players specifically between turns 25 and 35, we add the skill intervention vector to the model’s MLP bias term, which means that it is applied at every token in the input PGN string. We observe that it maintains effective performance across all stages of the game.
@@ 5 Related Work
There have been several examples of researchers finding semantics extraction in Large Language Models trained on text. In 2017, researchers trained a long short-term memory network (LSTM) to predict the next character in Amazon reviews and found that linear probes trained on the model exceeded the state of the art in sentiment classification (Radford et al., 2017). More recently, researchers have found evidence of concepts such as truthfulness and emotion in the Llama LLM (Li et al., 2023b; Zou et al., 2023).
@@ 6 Conclusion
Chess is a constrained setting compared to the domain of natural language. This constrained setting makes it possible to perform these nuanced measurements and interventions and provide valuable insights into the underlying mechanisms of steering vectors and other concepts in large language models. In addition, it also enables exploration and comparisons of a variety of interpretability and intervention methods. There is room for improvement in our intervention approaches, and we can easily measure our ability to make fine-grained changes to the model’s activations, such as by deleting a single piece from the board.
However, chess is ultimately a well-defined, rule-based system that does not capture the complexity and ambiguity present in less structured environments like natural language. Accurate understanding of LLM world models could enable many applications, such as detecting and reducing hallucinations. Our hope is that these techniques can become practical in these domains.
Acknowledgments
We thank Walt Woods, Alexander Grushin, Martin Wattenberg, Kenneth Li, Nikola Jurkovic, Jannik Brinkman, and Austin Davis for feedback and advice.
@@ Appendix C Intervention Technique Comparisons
We find that a single layer intervention can work almost as well as a multiple layer intervention. We had the following board intervention success rates per single layer intervention.
| Layer | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 |
@@ Appendix E Model evaluations against Stockfish
Appendix F Probe F1 Scores by Piece Type
| Piece Type | 8 Layer Model | 16 Layer Model | 8 Layer Random Init | 16 Layer Random Init |
| Opp. King | 0.985 | 0.983 | 0.658 | 0.667 |
| Opp. Queen | 0.986 | 0.986 | 0.573 | 0.576 |
| Opp. Rook | 0.973 | 0.987 | 0.638 | 0.647 |
| Opp. Bishop | 0.985 | 0.996 | 0.480 | 0.480 |
| Opp. Knight | 0.986 | 0.993 | 0.494 | 0.473 |
| Opp. Pawn | 0.980 | 0.997 | 0.604 | 0.602 |
| Blank | 0.993 | 0.997 | 0.821 | 0.818 |
| My Pawn | 0.997 | 0.994 | 0.641 | 0.631 |
| My Knight | 0.991 | 0.997 | 0.519 | 0.499 |
| My Bishop | 0.994 | 0.998 | 0.489 | 0.475 |
| My Rook | 0.986 | 0.995 | 0.635 | 0.646 |
| My Queen | 0.992 | 0.998 | 0.569 | 0.570 |
| My King | 0.998 | 0.997 | 0.684 | 0.644 |
Table 9 provides a detailed breakdown of the linear probe performance for each piece type and board state. The trained models (8-layer and 16-layer) show consistently high F1 scores across all piece types, with most scores above 0.98. In contrast, the randomly initialized models show much lower F1 scores, typically in the 0.5-0.6 range, with better performance for blank squares.